# Understanding Self-Admitted Technical Debt in Test Code: An Empirical Study

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Software design techniquesccs: Software and its engineering Documentationccs: Software and its engineering Software evolutionccs: Software and its engineering Maintaining softwareccs: Software and its engineering Empirical software validationccs: Software and its engineering Software reliability
@@ 2.1.3. Relation between SATD and software quality
Palomba et al. (Palomba et al., 2017) studied the relationship between SATD and refactoring in three OSS projects. They found that approximately 46% of refactored classes had SATD in previous versions. Furthermore, in 67% of cases, comments and descriptions related to SATD were deleted by refactoring, resolving the technical debt. This study shows that SATD might play an important role as a motivation for refactoring. Maldonado et al. (da S. Maldonado et al., 2017) analyzed the relationship between SATD and code smells. Code smells indicates potential flaws in the design or implementation of production code (Fowler, 1999), which are likely to cause problems in the future. The authors analyzed the overlap between SATD and three representative types of code smells: Long Method, God Class, and Feature Envy at the file level. Their investigation, which targeted 10 OSS repositories, revealed that 65% of files containing SATD also had Long Methods, 44.2% contained God Classes, and 20.7% contained Feature Envy classes. Moreover, when considering all code smells, they found that an average of 69.7% of files containing SATD had at least one code smell. From these findings, they concluded that while there are certain overlaps between SATD and code smells, these two indicators can serve as complementary approaches for detecting technical debt.
@@ 4.1. Data Collection
# Repository Description P.LOC T.LOC Files SATDs Stars Commits Forks Dev. Ref
1 Ant Project build tool 114,501 32,279 1,327 243 442 14,963 445 80 (da S. Maldonado et al., 2017)
2 ArgoUML UML modeling tool 160,769 16,705 1,908 1,828 262 17,797 102 5 (da S. Maldonado et al., 2017)
3 Columba E-mail client 66,394 1,526 993 138 - - - - (da S. Maldonado et al., 2017)
4 EMF Modeling tool 606,645 105,084 3,140 141 20 10,341 24 36 (da S. Maldonado et al., 2017)
5 Hibernate ORM FW 163,448 166,085 4,705 812 - - - - (da S. Maldonado et al., 2017)
6 Jedit Text editor 121,532 3,502 617 164 - - - - (da S. Maldonado et al., 2017)
7 JFreeChart Chart library 97,460 41,664 1,017 93 1,307 3,894 565 30 (da S. Maldonado et al., 2017)
8 Jmeter Perf. tester 120,620 28,143 1,402 272 8,921 18,254 2,193 80 (da S. Maldonado et al., 2017)
9 Jruby Ruby interpreter 265,506 10,427 1,793 1,016 3,835 53,597 928 362 (da S. Maldonado et al., 2017)
10 Squirrel SQL client 186,681 28,553 2,325 253 75 7,766 19 20 (da S. Maldonado et al., 2017)
11 Dubbo RPC FW 154,678 108,908 3,581 168 41,184 7,535 26,527 401 (Guo et al., 2019)
12 Gradle Project build tool 494,771 17,712 10,392 698 17,886 114,478 4,984 345 (Guo et al., 2019)
13 Groovy Script language 191,447 12,262 1,793 462 5,336 20,871 1,911 359 (Guo et al., 2019)
14 Hive Data warehouse 867,928 178,970 5,086 1,354 5,747 10,189 4,757 257 (Guo et al., 2019)
15 Maven Project build tool 55,245 19,455 1,012 173 4,716 10,966 2,764 209 (Guo et al., 2019)
16 Poi File handler 281,995 142,878 3,710 818 2,062 12,828 792 17 (Guo et al., 2019)
17 Spring Framework Java App FW 383,270 391,933 8,310 283 58,540 29,550 38,600 355 (Guo et al., 2019)
18 Storm Computation system 249,986 40,390 2,145 103 6,644 10,798 4,064 281 (Guo et al., 2019)
19 Tomcat Web server 268,488 92,832 2,616 814 7,891 25,797 5,211 133 (Guo et al., 2019)
20 Zookeeper Coordination service 61,382 59,789 925 109 12,549 2,534 7,295 238 (Guo et al., 2019)
21 Commons IO IO library 17,890 33,866 513 87 1,036 4,849 695 115 (Kashiwa et al., 2021)
22 Spring SQL mapping FW 2,079 4,240 123 5 2,886 2,165 2,628 49 (Kashiwa et al., 2021)
23 Joda-Beans Code generator 17,353 34,625 328 2 146 891 40 11 (Kashiwa et al., 2021)
24 Jsoup HTML library 15,306 14,745 153 58 11,222 1,969 2,248 103 (Kashiwa et al., 2021)
25 Spark Web FW 6,222 5,048 184 5 9,660 1,041 1,569 123 (Kashiwa et al., 2021)
26 LittleProxy HTTP proxy 4,180 4,665 88 3 2,095 998 782 27 (Kashiwa et al., 2021)
27 RxJava JDBC Database library 4,611 3,340 79 7 803 926 115 13 (Kashiwa et al., 2021)
28 Spoon Program analysis library 83,543 83,893 2,407 148 1,840 4,651 361 134 (Kashiwa et al., 2021)
29 Accumulo Key-value store 441,912 44,357 2,179 110 1,104 9,001 463 158 (Martins et al., 2023)
30 BookKeeper Storage service 157,229 121,242 2,351 62 1,955 3,322 960 204 (Martins et al., 2023)
31 Camel Integration FW 1,041,075 653,891 21,464 1,999 5,906 70,277 5,048 331 (Martins et al., 2023)
32 Cassandra NoSQL database 314,778 273,241 3,749 496 9,292 28,567 3,723 272 (Martins et al., 2023)
33 CXF Web services FW 389,365 308,696 7,513 443 885 18,004 1,442 238 (Martins et al., 2023)
34 Flink Processing engine 838,513 748,068 13,521 696 25,086 34,899 13,717 285 (Martins et al., 2023)
35 Hadoop Processing FW 997,669 902,657 11,871 1,022 15,190 27,121 9,076 303 (Martins et al., 2023)
36 Kafka Event streaming platform 271,653 356,138 4,245 327 30,571 11,728 14,524 347 (Martins et al., 2023)
37 Karaf Modulith runtime 105,567 23,311 1,603 135 691 9,471 664 163 (Martins et al., 2023)
38 Wicket App FW 142,446 77,054 3,306 141 768 21,683 393 116 (Martins et al., 2023)
39 Struts App FW 123,978 87,049 2,589 179 1,318 7,160 824 77 (Bui et al., 2022)
40 Compress Archiving library 43,544 29,119 623 170 370 5,121 295 95 (Bui et al., 2022)
41 Jenkins CI/CD server 113,619 75,714 1,782 509 24,213 34,958 9,093 296 (Bui et al., 2022)
42 Spring Security Auth. FW 103,939 167,283 3,145 64 9,224 15,713 6,092 388 (Bui et al., 2022)
43 FileUpload File upload library 2,221 4,205 92 13 248 1,657 187 49 (Bui et al., 2022)
44 Imaging Image library 30,710 11,347 623 73 461 2,486 197 45 (Bui et al., 2022)
45 Sling Web FW 7,246 5,750 179 8 15 687 22 21 (Bui et al., 2022)
46 UAA Auth. FW 62,053 124,466 1,646 24 1,625 10,952 834 158 (Bui et al., 2022)
47 Jackson XML data binding 6,283 10,702 195 9 600 1,560 230 41 (Bui et al., 2022)
48 Prime JWT Token management 3,759 2,809 101 4 187 326 43 12 (Bui et al., 2022)
49 OpenRefine Data cleaning tool 57,094 28,418 1,026 213 11,446 7,995 2,080 353 (Bui et al., 2022)
50 Quarkus App FW 540,637 449,513 16,729 812 14,764 45,217 2,900 391 (Soares et al., 2023)
@@
For each of these repositories, we clone and checkout the latest revision of the master/main branch. However, for some repositories, we collected specific snapshot versions instead. We then identify SATD in the production code and the test code. Since all 50 repositories in our study are Java projects using either Maven or Gradle, we identified production and test directories by analyzing their build configuration files. We located each repository’s pom.xml file (Maven projects) or build.gradle file (Gradle projects) and parsed it to extract directory paths. For Maven projects, we extracted paths from the sourceDirectory and testSourceDirectory elements; for Gradle projects, we identified paths from the main and test properties. When these paths were not explicitly defined, we used standard Java conventions: src/main/java for production code and src/test/java for test code [^5] https://maven.apache.org/guides/introduction/introduction-to-the-standard-directory-layout.html [^6] https://docs.gradle.org/current/userguide/java_plugin.html#sec:java_project_layout. In the end, we identify 100,364 files in the production directories and 60,515 files in the test directories. From these files, we identify SATDs in the production code and test code. We consider SATD instances found in the production code as Production SATDs and those found in the test code as Test SATDs. For SATD detection, we employ a state-of-the-art SATD detection tool, DebtHunter (Sala et al., 2021). DebtHunter is an ML-based tool that uses the Sequential Minimal Optimization (SMO) algorithm (Chawla et al., 2002). A previous study (Sala et al., 2021) has demonstrated the high performance of DebtHunter, with a precision of 0.972, recall of 0.967, and F1-score of 0.965, outperforming other SATD detection tools. Additionally, since DebtHunter has been trained with multiple labels for classifying the types of SATD, we also use this tool to identify SATD types. Note that we modified the source code of DebtHunter so that it can detect Test SATD because the original version excludes the test code.
@@ 4.2.2. RQ2 (Relationship with Test Quality)
The test smells are detected with tsDetect (Peruma et al., 2020) [^8] https://github.com/bin-lin/TestSmellDetector, a state-of-the-art tool which can detect 16 of 19 test smells proposed by van Deursen et al. (van Deursen et al., 2001) at the method level. Their evaluation shows an average precision and recall exceeding 95% for detecting each type of smell. To detect code smells, we use Designite Java [^9] https://github.com/tushartushar/DesigniteJava, which can detect 10 of 11 code smells proposed by Sharma et al. (Sharma et al., 2017). To ensure the reliability of these automated detectors on our specific dataset, we conducted an experimental manual validation. We randomly sampled 100 instances identified as test smells by tsDetect and 100 instances identified as code smells by Designite Java. Two of the authors then independently inspected each instance to verify if it was a true positive. Any disagreements were resolved through discussion with a third inspector to reach a consensus. Our validation confirmed a high precision for both tools: 90% for the test smell detection and 94% for the code smell detection. These results to a high extent guarantee the validity of the quality metrics used in our analysis.
@@ 5.1. $RQ_{1}$: How prevalent is SATD in test code?
As the lines of code in the production code and test code differ significantly, we also compare the normalized number of SATD instances. Figure 3 depicts the distribution of the numbers of SATD per 10,000 lines of code for each repository. In these violin plots, the thickness of the outer layer represents the probability density of the plotted values. In the center of each violin plot, the yellow line shows the median while the box bar represents the interquartile range. When looking at the quartiles, the numbers of SATD instances per 10k LOC range from 7.3 to 18.9 in the production code, while from 1.6 to 14.1 in the test code. The average and median values for Production SATD is also much higher than that for Test SATD (average: 15.6 vs. 9.5, median: 10.9 vs. 4.1). In terms of median, the normalized number of Test SATD is 62.4% smaller than that of the Production SATDs. A statistically significant difference is also observed with the Wilcoxon signed-rank test (p-value: 0.00006). In addition, the effect size (Cliff’s Delta) was found to be 0.365, indicating a medium difference. Table 2 presents the detailed results of the statistical test.
@@ 5.2. $RQ_{2}$: Is SATD in test code correlated with test quality?
Furthermore, we conducted a correlation analysis among these quality metrics. We calculated Spearman’s rank correlation coefficient for all pairs of the seven metrics, separately for test methods with and without SATD. The Spearman coefficient ($\rho$) ranges from -1 to +1, indicating the direction and strength of a monotonic relationship. A positive value indicates that as one metric increases, the other tends to increase, while a negative value indicates that as one metric increases, the other tends to decrease. Figure 7 and 7 visualize these correlation coefficients as two heatmaps.
@@
Figure 6. Spearman Correlation Matrix (With SATD)
Figure 7. Spearman Correlation Matrix (Without SATD)
@@ 5.3. $RQ_{3}$: What are the purposes of SATD in test code?
Main Cat. Sub Cat. Common words Detailed Category      #
Production- Failures { fails, test, enabled, passes, macos } Indicates the failure-triggering environments/settings 11
@@
issues On-hold { currently, supported, yet, work, fix } Indicates the existence of defects in production code 42
@@
Test- Test { assert, value, test, check, exception } Indicates incomplete or unimplemented tests 120
@@
issues Test Design { test, workaround, need, way, better } Workaround for impossible/difficult implementations 53
@@
Test { remove, version, buffer, response, future } Asks for test updates in response to future/latest software versions 17
@@
Of the 506 instances, 407 instances (80.4%) are assigned with consistent labels from both reviewers, resulting in a Cohen’s kappa coefficient of 0.78, indicating a substantial agreement according to Landis and Koch (Landis and Koch, 1977). The conflicts were resolved with a third annotator. Throughout the manual inspection, we found that 36 instances were false positive (incorrectly detected as SATD) and the label of 40 instances could not be assigned due to the lack of information. As a result, 430 SATD instances were included in the final results. Table 5 shows the final classification of Test SATD and the breakdown of each category. In Table 5, we identified the top five most frequent words for each category except typical SATD tags such as TODO and FIXME. In particular, we split the words in SATD comments, removed stopwords using the NLTK library, [^10] https://www.nltk.org and identified the most frequent words for each category.
In the following, we introduce each category, separated by the origin of issues (whether the issue originated from production-code or test code).

### v2

@@ Abstract.
journal: TOSEMcopyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Software design techniquesccs: Software and its engineering Documentationccs: Software and its engineering Software evolutionccs: Software and its engineering Maintaining softwareccs: Software and its engineering Empirical software validationccs: Software and its engineering Software reliability
This manuscript is accepted at ACM Transactions on Software Engineering and Methodology (TOSEM).
@@ 2.1.3. Relation between SATD and software quality
Palomba et al. (Palomba et al., 2017) studied the relationship between SATD and refactoring in three OSS projects. They found that approximately 46% of refactored classes had SATD in previous versions. Furthermore, in 67% of cases, comments and descriptions related to SATD were deleted by refactoring, resolving the technical debt. This study shows that SATD might play an important role as a motivation for refactoring. Maldonado et al. (da S. Maldonado et al., 2017) analyzed the relationship between SATD and code smells. Code smells indicate potential flaws in the design or implementation of production code (Fowler, 1999), which are likely to cause problems in the future. The authors analyzed the overlap between SATD and three representative types of code smells: Long Method, God Class, and Feature Envy at the file level. Their investigation, which targeted 10 OSS repositories, revealed that 65% of files containing SATD also had Long Methods, 44.2% contained God Classes, and 20.7% contained Feature Envy classes. Moreover, when considering all code smells, they found that an average of 69.7% of files containing SATD had at least one code smell. From these findings, they concluded that while there are certain overlaps between SATD and code smells, these two indicators can serve as complementary approaches for detecting technical debt.
@@ 4.1. Data Collection
# Repository P.LOC T.LOC Files SATDs Stars Commits Forks Dev. Ref
1 Ant 114,501 32,279 1,327 243 442 14,963 445 80 (da S. Maldonado et al., 2017)
2 ArgoUML 160,769 16,705 1,908 1,828 262 17,797 102 5 (da S. Maldonado et al., 2017)
3 Columba 66,394 1,526 993 138 - - - - (da S. Maldonado et al., 2017)
4 EMF 606,645 105,084 3,140 141 20 10,341 24 36 (da S. Maldonado et al., 2017)
5 Hibernate 163,448 166,085 4,705 812 - - - - (da S. Maldonado et al., 2017)
6 Jedit 121,532 3,502 617 164 - - - - (da S. Maldonado et al., 2017)
7 JFreeChart 97,460 41,664 1,017 93 1,307 3,894 565 30 (da S. Maldonado et al., 2017)
8 Jmeter 120,620 28,143 1,402 272 8,921 18,254 2,193 80 (da S. Maldonado et al., 2017)
9 Jruby 265,506 10,427 1,793 1,016 3,835 53,597 928 362 (da S. Maldonado et al., 2017)
10 Squirrel 186,681 28,553 2,325 253 75 7,766 19 20 (da S. Maldonado et al., 2017)
11 Dubbo 154,678 108,908 3,581 168 41,184 7,535 26,527 401 (Guo et al., 2019)
12 Gradle 494,771 17,712 10,392 698 17,886 114,478 4,984 345 (Guo et al., 2019)
13 Groovy 191,447 12,262 1,793 462 5,336 20,871 1,911 359 (Guo et al., 2019)
14 Hive 867,928 178,970 5,086 1,354 5,747 10,189 4,757 257 (Guo et al., 2019)
15 Maven 55,245 19,455 1,012 173 4,716 10,966 2,764 209 (Guo et al., 2019)
16 Poi 281,995 142,878 3,710 818 2,062 12,828 792 17 (Guo et al., 2019)
17 Spring Framework 383,270 391,933 8,310 283 58,540 29,550 38,600 355 (Guo et al., 2019)
18 Storm 249,986 40,390 2,145 103 6,644 10,798 4,064 281 (Guo et al., 2019)
19 Tomcat 268,488 92,832 2,616 814 7,891 25,797 5,211 133 (Guo et al., 2019)
20 Zookeeper 61,382 59,789 925 109 12,549 2,534 7,295 238 (Guo et al., 2019)
21 Commons IO 17,890 33,866 513 87 1,036 4,849 695 115 (Kashiwa et al., 2021)
22 Spring 2,079 4,240 123 5 2,886 2,165 2,628 49 (Kashiwa et al., 2021)
23 Joda-Beans 17,353 34,625 328 2 146 891 40 11 (Kashiwa et al., 2021)
24 Jsoup 15,306 14,745 153 58 11,222 1,969 2,248 103 (Kashiwa et al., 2021)
25 Spark 6,222 5,048 184 5 9,660 1,041 1,569 123 (Kashiwa et al., 2021)
26 LittleProxy 4,180 4,665 88 3 2,095 998 782 27 (Kashiwa et al., 2021)
27 RxJava JDBC 4,611 3,340 79 7 803 926 115 13 (Kashiwa et al., 2021)
28 Spoon 83,543 83,893 2,407 148 1,840 4,651 361 134 (Kashiwa et al., 2021)
29 Accumulo 441,912 44,357 2,179 110 1,104 9,001 463 158 (Martins et al., 2023)
30 BookKeeper 157,229 121,242 2,351 62 1,955 3,322 960 204 (Martins et al., 2023)
31 Camel 1,041,075 653,891 21,464 1,999 5,906 70,277 5,048 331 (Martins et al., 2023)
32 Cassandra 314,778 273,241 3,749 496 9,292 28,567 3,723 272 (Martins et al., 2023)
33 CXF 389,365 308,696 7,513 443 885 18,004 1,442 238 (Martins et al., 2023)
34 Flink 838,513 748,068 13,521 696 25,086 34,899 13,717 285 (Martins et al., 2023)
35 Hadoop 997,669 902,657 11,871 1,022 15,190 27,121 9,076 303 (Martins et al., 2023)
36 Kafka 271,653 356,138 4,245 327 30,571 11,728 14,524 347 (Martins et al., 2023)
37 Karaf 105,567 23,311 1,603 135 691 9,471 664 163 (Martins et al., 2023)
38 Wicket 142,446 77,054 3,306 141 768 21,683 393 116 (Martins et al., 2023)
39 Struts 123,978 87,049 2,589 179 1,318 7,160 824 77 (Bui et al., 2022)
40 Compress 43,544 29,119 623 170 370 5,121 295 95 (Bui et al., 2022)
41 Jenkins 113,619 75,714 1,782 509 24,213 34,958 9,093 296 (Bui et al., 2022)
42 Spring Security 103,939 167,283 3,145 64 9,224 15,713 6,092 388 (Bui et al., 2022)
43 FileUpload 2,221 4,205 92 13 248 1,657 187 49 (Bui et al., 2022)
44 Imaging 30,710 11,347 623 73 461 2,486 197 45 (Bui et al., 2022)
45 Sling 7,246 5,750 179 8 15 687 22 21 (Bui et al., 2022)
46 UAA 62,053 124,466 1,646 24 1,625 10,952 834 158 (Bui et al., 2022)
47 Jackson 6,283 10,702 195 9 600 1,560 230 41 (Bui et al., 2022)
48 Prime JWT 3,759 2,809 101 4 187 326 43 12 (Bui et al., 2022)
49 OpenRefine 57,094 28,418 1,026 213 11,446 7,995 2,080 353 (Bui et al., 2022)
50 Quarkus 540,637 449,513 16,729 812 14,764 45,217 2,900 391 (Soares et al., 2023)
@@
For each of these repositories, we clone and checkout the latest revision of the master/main branch. However, for some repositories, we collected specific snapshot versions instead. We then identify SATD in the production code and the test code. Since all 50 repositories in our study are Java projects using either Maven or Gradle, we identified production and test directories by analyzing their build configuration files. We located each repository’s pom.xml file (Maven projects) or build.gradle file (Gradle projects) and parsed it to extract directory paths. For Maven projects, we extracted paths from the sourceDirectory and testSourceDirectory elements; for Gradle projects, we identified paths from the main and test properties. When these paths were not explicitly defined, we used standard Java conventions: src/main/java for production code and src/test/java for test code. [^5] https://maven.apache.org/guides/introduction/introduction-to-the-standard-directory-layout.html, [^6] https://docs.gradle.org/current/userguide/java_plugin.html#sec:java_project_layout In the end, we identify 100,364 files in the production directories and 60,515 files in the test directories. From these files, we identify SATDs in the production code and test code. We consider SATD instances found in the production code as Production SATDs and those found in the test code as Test SATDs. For SATD detection, we employ a state-of-the-art SATD detection tool, DebtHunter (Sala et al., 2021). DebtHunter is an ML-based tool that uses the Sequential Minimal Optimization (SMO) algorithm (Chawla et al., 2002). A previous study (Sala et al., 2021) has demonstrated the high performance of DebtHunter, with a precision of 0.972, recall of 0.967, and F1-score of 0.965, outperforming other SATD detection tools. Additionally, since DebtHunter has been trained with multiple labels for classifying the types of SATD, we also use this tool to identify SATD types. Note that we modified the source code of DebtHunter so that it can detect Test SATD because the original version excludes the test code.
@@ 4.2.2. RQ2 (Relationship with Test Quality)
The test smells are detected with tsDetect (Peruma et al., 2020), [^8] https://github.com/bin-lin/TestSmellDetector a state-of-the-art tool which can detect 16 of 19 test smells proposed by van Deursen et al. (van Deursen et al., 2001) at the method level. Their evaluation shows an average precision and recall exceeding 95% for detecting each type of smell. To detect code smells, we use Designite Java, [^9] https://github.com/tushartushar/DesigniteJava which can detect 10 of 11 code smells proposed by Sharma et al. (Sharma et al., 2017). To ensure the reliability of these automated detectors on our specific dataset, we conducted an experimental manual validation. We randomly sampled 100 instances identified as test smells by tsDetect and 100 instances identified as code smells by Designite Java. Two of the authors then independently inspected each instance to verify if it was a true positive. Any disagreements were resolved through discussion with a third inspector to reach a consensus. Our validation confirmed a high precision for both tools: 90% for the test smell detection and 94% for the code smell detection. These results to a high extent guarantee the validity of the quality metrics used in our analysis.
@@ 5.1. $RQ_{1}$: How prevalent is SATD in test code?
As the lines of code in the production code and test code differ significantly, we also compare the normalized number of SATD instances. Figure 3 depicts the distribution of the numbers of SATD per 10,000 lines of code for each repository. In these violin plots, the thickness of the outer layer represents the probability density of the plotted values. In the center of each violin plot, the yellow line shows the median while the box bar represents the interquartile range. When looking at the quartiles, the number of SATD instances per 10k LOC range from 7.3 to 18.9 in the production code, while from 1.6 to 14.1 in the test code. The average and median values for Production SATD are also much higher than that for Test SATD (average: 15.6 vs. 9.5, median: 10.9 vs. 4.1). In terms of median, the normalized number of Test SATD is 62.4% smaller than that of the Production SATDs. A statistically significant difference is also observed with the Wilcoxon signed-rank test (p-value: 0.00006). In addition, the effect size (Cliff’s Delta) was found to be 0.365, indicating a medium difference. Table 2 presents the detailed results of the statistical test.
@@ 5.2. $RQ_{2}$: Is SATD in test code correlated with test quality?
Furthermore, we conducted a correlation analysis among these quality metrics. We calculated Spearman’s rank correlation coefficient for all pairs of the seven metrics, separately for test methods with and without SATD. The Spearman coefficient ($\rho$) ranges from -1 to +1, indicating the direction and strength of a monotonic relationship. A positive value indicates that as one metric increases, the other tends to increase, while a negative value indicates that as one metric increases, the other tends to decrease. Figure 6(a) and 6(b) visualize these correlation coefficients as two heatmaps.
@@
(a) With SATD
(b) Without SATD
Figure 6. Spearman Correlation Matrices with and without SATD
@@ 5.3. $RQ_{3}$: What are the purposes of SATD in test code?
Main Cat. Sub Cat. Detailed Category      #
Production- Failures Indicates the failure-triggering environments/settings 11
@@
issues On-hold Indicates the existence of defects in production code 42
@@
Test- Test Indicates incomplete or unimplemented tests 120
@@
issues Test Design Workaround for impossible/difficult implementations 53
@@
Test Asks for test updates in response to future/latest software versions 17
@@
Of the 506 instances, 407 instances (80.4%) are assigned with consistent labels from both reviewers, resulting in a Cohen’s kappa coefficient of 0.78, indicating a substantial agreement according to Landis and Koch (Landis and Koch, 1977). The conflicts were resolved with a third annotator. Throughout the manual inspection, we found that 36 instances were false positives (incorrectly detected as SATD) and the label of 40 instances could not be assigned due to the lack of information. As a result, 430 SATD instances were included in the final results. Table 5 shows the final classification of Test SATD and the breakdown of each category. In Table 6, we identified the top five most frequent words for each category except typical SATD tags such as TODO and FIXME. In particular, we split the words in SATD comments, removed stopwords using the NLTK library, [^10] https://www.nltk.org and identified the most frequent words for each category.
Table 6. Most frequent words by subcategory
Subcategory Common words
Failures fails, test, enabled, passes, macos
On-hold Tasks currently, supported, yet, work, fix
Test Completeness assert, value, test, check, exception
Test Design and Implementation test, workaround, need, way, better
Test Maintenance remove, version, buffer, response, future
In the following, we introduce each category, separated by the origin of issues (whether the issue originated from production code or test code).